The distance between Kansas City, Missouri and Denver is 600 miles. On a certain wall map, this is represented by a length of 2.4 ft. On the map, how many feet would there be between Memphis and Philadelphia, two cities that are actually 1000 miles apart?

Answer by checkley77(12844)

2.4/600=x/1000
600x=2,400
x=2,400/600
x=4 ft. between Memphis & Philadelphia on the map.